OpenStudy (anonymous):

solve 5*3^x=5, explain please !: )

OpenStudy (anonymous):

\[ 5\times3^x=5\\ 3^x=1\\ x=\log_3 1=0 \]

OpenStudy (anonymous):

5*3^x=5 3^x =1 or 3^x=3^0 hence x=0